OpenStudy (anonymous):

\[sin(2~arc~cos{3\over5})\]i forgot how to do this...

OpenStudy (anonymous):

@artofspeed

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

sin(2x)=2 * sinx * cos x

OpenStudy (anonymous):

arccos(3/5) = x can be drawn as

OpenStudy (anonymous):

hence the third side is sqrt(5^2-3^2)=4

OpenStudy (anonymous):

sinx = 4/5 and cosx=3/5

OpenStudy (anonymous):

sin(2x)=2 * sinx * cos x = 2*4/5*3/5 = 24/25

OpenStudy (anonymous):

i dont get how sin2x=2sinxcosx

OpenStudy (anonymous):

nevermind...

OpenStudy (anonymous):

thanks @artofspeed ! :)

OpenStudy (anonymous):

nooo problem

OpenStudy (anonymous):

lol :D